1. Hydrogen and oxygen react chemically to form water. How much water would form if 19.2 grams of hydrogen reacted with 40.1 grams of oxygen? (H2 + O2 → H2O) 
1
Expert's answer
2020-10-04T09:20:17-0400

Balanced reaction:

2H2 + O2 → 2H2O

M(H2) = 2 g/mol

M(O2) = 32 g/mol

"n = \frac{m}{M}"

n(H2) "= \frac{19.2}{2} = 9.6 \;mol"

n(O2) "= \frac{40.1}{32} = 1.25 \;mol"

Proportion from the reaction:

H2: O2 = 2:1

Proportion from the reactants masses:

H2: O2 = 9.6:1.25

O2 is the limiting reactant that will completely react.

n(H2O) = 2n(O2) "= 2\times1.25 = 2.5\; mol"

M(H2O) = 18 g/mol

m(H2O) "= n\times M = 2.5\; mol\times 18\;g/mol = 45\; g"

Answer: 45 g